### v1

# Domain Generalization: A Tale of Two ERMs

Abstract
Domain generalization (DG) is the problem of generalizing from several distributions (or domains), for which labeled training data are available, to a new test domain for which no labeled data is available. A common finding in the DG literature is that it is difficult to outperform empirical risk minimization (ERM) on the pooled training data.
In this work, we argue that this finding has primarily been reported for datasets satisfying a covariate shift assumption. When the dataset satisfies a posterior drift assumption instead, we show that “domain-informed ERM,” wherein feature vectors are augmented with domain-specific information, outperforms pooling ERM. These claims are supported by a theoretical framework and experiments on language and vision tasks.
fnum@section1 Introduction:
the ERM dilemma
Domain generalization (DG) is a learning problem where the learner has access to labeled data from several source domains, and the goal is to generalize to a new target domain for which no labeled data is available. Let $X$ denote the input features, $Y$ the label, and $D$ the domain index.
A persistent puzzle in DG is the surprising effectiveness of empirical risk minimization (ERM), a baseline that simply pools labeled data from all source domains together and trains a domain-agnostic classifier. Despite extensive efforts to design sophisticated DG algorithms, multiple studies have consistently shown that ERM remains highly competitive:
  • Gulrajani and Lopez-Paz (2021) (empirical): “when carefully implemented and tuned, ERM outperforms the state-of-the-art in terms of average performance… no algorithm included in DomainBed (dataset) outperforms ERM by more than $1\%$.”
  • Rosenfeld et al. (2021) (theory): “Invariant Risk Minimization and its alternatives fundamentally do not improve over standard ERM.”
  • Teterwak et al. (2025) (empirical): “the additional tuning in our improved baseline ERM++ outperforms both the prior ERM baselines and all recent SOTA methods on DomainBed.”
Similar observations about the strong performance of ERM have been made on other DG benchmarks (Koh et al., 2021; Sagawa et al., 2022), as well as in related settings such as federated domain generalization (Bai et al., 2024).
A related observation is that most existing DG approaches learn a classifier that predicts $Y$ solely from $X$, thus ignoring domain information during inference. This is reflected in recent surveys:
  • Wang et al. (2022) (survey): “The goal of domain generalization is to learn a robust and generalizable predictive function $h:\mathcal{X}\to\mathcal{Y}$ from the $M$ training domains to achieve a minimum prediction error on an unseen test domain $S_{\text{test}}$.”
  • Zhou et al. (2023) (survey): “The goal of DG is to learn a predictive model $f:\mathcal{X}\to\mathcal{Y}$ using only source domain data such that the prediction error on an unseen domain $T=\left\{x^{T}\right\}$ is minimized.”
This is despite the fact that early works on DG learn predictions based not only on feature vectors, but also on domain-specific information (Blanchard et al., 2011; Muandet et al., 2013). While more recent methods also incorporate domain information into the training and inference procedure (Dubey et al., 2021; Bui et al., 2021; Yao et al., 2024), the fundamental role of domain information—its precise decision-theoretic value and the conditions under which it strictly improves performance—remains poorly understood.
In this work, we argue that the prevailing conclusions about ERM being “hard to beat” stem primarily from the fact that most benchmark DG datasets are from vision tasks. These datasets are characterized by a covariate shift assumption, which means that there exists a single classifier that performs well on all domains, and only the marginal distribution of $X$ changes from domain to domain. In such applications, strong performance is indeed possible without the use of domain-specific information, and ERM on pooled training data can perform quite well.
We consider DG problems characterized by posterior drift, where the conditional distribution of $Y|X$ (i.e., the posterior) changes with domain. We argue that for such DG problems, pooling ERM is inadequate, and stronger performance is achievable by “domain-informed” ERM, where domain-specific information is used both during training and at inference.
More broadly, a key goal of this work is to reassert (and extend) the importance of the original formulation of DG, which leverages domain-specific information when making predictions. We argue that it is this element of DG that makes it fundamentally distinct from conventional classification.
The contributions of this work are:
  • A theoretical framework extending the original formulation of DG by Blanchard et al. (2011).
  • Risk bounds that characterize when domain-specific information is beneficial (posterior drift) and when it is not (covariate shift).
  • A quantification of the difference between DG and domain adaptation, addressing an open question in Blanchard et al. (2021, Lemma 9).
  • Empirical validation of these findings on both language and vision tasks.
fnum@section2 Literature review
Blanchard et al. (2011) introduced domain generalization (DG), motivated by a medical application involving the automatic gating of flow cytometry data. Since then, most DG research has focused on applications in computer vision. A typical DG task in this setting involves training models on labeled images from multiple visual domains (e.g., styles or rendering conditions) and evaluating generalization to a previously unseen domain. Benchmark datasets such as VLCS (Fang et al., 2013), PACS (Li et al., 2017), OfficeHome (Venkateswara et al., 2017), DomainNet (Peng et al., 2019), and ImageNet-Sketch (Wang et al., 2019) have become standard in this line of work.
In these vision-based setups, the underlying distributional shift can be described as covariate shift (Ben-David et al., 2006; Mansour et al., 2009), where the marginal distribution $P_{X}$ varies significantly across domains, while the posterior $P_{Y|X}$ is constant. Importantly, domain information is frequently viewed as irrelevant or even spurious (Sagawa et al., 2020) for predicting labels. Consequently, much of the literature has focused on learning domain-invariant representations (Sun and Saenko, 2016; Ganin et al., 2016; Arjovsky et al., 2019). Additional references are in Appendix˜A.
In contrast, our work is motivated by a different class of problems characterized by posterior drift (Scott, 2019; Cai and Wei, 2021; Maity et al., 2024; Zhu et al., 2024; Wang and Ning, 2025), where the conditional distribution of $Y|X$ varies across domains. This type of shift commonly arises in natural language processing (NLP). For a given sentence $X$, different annotators—or populations—may interpret its semantic content differently, leading to divergent labels $Y$ (e.g., offensive vs. non-offensive, positive vs. negative). Such inherent ambiguity in language often results in systematic disagreement in annotations. These effects have been documented across a wide range of NLP tasks (De Marneffe et al., 2019; Plank, 2022; Deng et al., 2023).
fnum@section3 A general framework
In standard classification, a random pair $(X,Y)$ is assumed to be drawn from a fixed joint distribution $P_{XY}$, where $X\in\mathcal{X}$ is a feature vector and $Y\in\mathcal{Y}=\{1,\ldots,K\}$ the corresponding class label [^1] This section easily extends to regression, but subsequent sections are specific to classification.. The goal is to learn a function $f:\mathcal{X}\to\mathcal{Y}$ that minimizes the risk
$\mathbb{E}_{(X,Y)\sim P_{XY}}\left[\mathbbm{1}_{f(X)\neq Y}\right].$
Domain generalization (DG) can be framed in a similar way. Let $\mathcal{D}$ denote a set of possible domains, where the term domain is a synonym for a joint distribution of $X$ and $Y$. Let $D$ be a random variable on $\mathcal{D}$. Furthermore, let $M$ be a random variable on a space $\mathcal{M}$ that, intuitively, provides partial information about $D$. The idea in DG is that $D$ determines a distribution of $(X,Y)$, but is not observed. $M$ provides partial information (or metadata) about $D$, and is thus useful at test time in adapting the classifier to the test domain.
Remark 1
While the choice of $M$ will depend on the application, one choice that is always viable is to take $M$ to be $P_{X|D}$, the marginal distribution $X$ for the given domain, which is known at test time though the unlabeled test sample.
As we argue below, the observability of $M$ is what makes DG an interesting problem, and distinct from standard classification.
Formally, we assume that $(X,Y,M,D)$ are jointly distributed, with joint distribution denoted $P_{XYMD}$. This distribution induces several other distributions of interest in this paper. We follow convention in denoting marginal distributions by keeping the relevant subscripts. For example, $P_{XYD}$ denotes the joint distribution of $(X,Y,D)$ after $M$ is marginalized out. Similarly, $P_{XY}$ denotes the marginal distribution of $(X,Y)$.
For any fixed $d\in\mathcal{D}$, $P_{XY|D=d}$ is a joint distribution of $(X,Y)$. Note that our notation is somewhat redundant, as both $d$ and $P_{XY|D=d}$ are notations for a domain == a joint distribution of $(X,Y)$, but these two notations both play a role our discussion. [^2] Blanchard et al. (2011, 2021) use $P_{XY}$ to denote a random domain, whereas in our notation, a random domain is either $P_{XY|D}$ or just $D$. Our introduction of $D$ for a random domain allows us to use $P_{XY}$ for the “average” domain, which will be a critical concept in what follows.
To formalize the notion that $M$ is a partial summary of $D$, we assume that $(X,Y)$ and $M$ are conditionally independent, given $D$:
$P_{XY|D,M}=P_{XY|D}.$ (1)
This implies that, given $D$, the joint distribution of $X$ and $Y$ does not change with knowledge of $M$. An important special case where this holds is when $M=g(D)$ for some deterministic $g:\mathcal{D}\to\mathcal{M}$. Table 1 lists three applications that illustrate this probabilistic framework, and that will be developed in Section˜6.
Table 1: Examples of domains and metadata in different tasks.
Task Input $X$ Label $Y$ Domain $D$ Metadata $M$
Sentiment analysis
(Multiple Annotators)
Sentence
Sentiment label
(e.g., positive, negative)
Annotator identity
(e.g., “Annotator 1”)
Annotator’s demographic profile
(e.g., age)
Review rating prediction
(Multiple Reviewers)
Product review
Numerical rating
(e.g., 1–5 stars)
Reviewer identity
(e.g., “Reviewer 2”)
Unlabeled texts written by the reviewer
$\{X_{i}\}_{i=1}^{n}\stackrel{{\scriptstyle iid}}{{\sim}}P_{X|D=d}$
Image classification
across styles
Image
Object category label
(e.g., dog, car)
Image style
(e.g., photograph, painting)
Textual description of style
The training data available to the learner is generated as follows: First, $N$ domains $d_{1},\ldots,d_{N}$ are sampled iid from $P_{D}$, but not observed. Then, conditioned on these $d_{i}$, corresponding values $m_{i}$ are observed. In addition, for each $i$, $1\leq i\leq N$, data $(x_{ij},y_{ij})$ are sampled iid from $P_{XY|D=d_{i}}$, $1\leq j\leq n_{i}$. In summary, the overall training data is
$\left(m_{i},(x_{ij},y_{ij})_{j=1}^{n_{i}}\right)_{i=1}^{N}.$
The goal of the learner is to produce a function $f$ that accurately predicts labels on a new, random domain. In particular, $f$ should minimize the risk
$R(f):=\mathbb{E}_{X,Y,M,D}\left[\mathbbm{1}_{f(\cdot)\neq Y}\right].$
In practice, this risk is estimated by holding out several of the domains, and averaging the test errors on them.
Remark 2
This probabilistic framing of DG generalizes the original formulation of DG by Blanchard et al. (2011). They focus on the special case where $M$ is the marginal distribution of $X$ for the given domain (aka, $M=P_{X|D}$), and focus on the challenges associated to learning from empirical samples of the training and test $X$-marginals.
The training setup described above naturally gives rise to two different ways of using the available data. On one hand, the learner may choose to ignore the domain information and simply pool together all training samples, treating them as if they were drawn iid from a single domain. On the other hand, the learner may choose to leverage the observed metadata $m_{i}$, which serves as side information about the underlying domain. These two strategies lead to two corresponding empirical risk minimization principles. Thus, let $\mathcal{F}\subset\{\mathcal{X}\times\mathcal{M}\to\mathcal{Y}\}$ denote a class of functions that take both instance $x$ and metadata $m$ as input, and $\mathcal{G}\subset\{\mathcal{X}\to\mathcal{Y}\}$ a class of functions that take only $x$ as input. Consider two empirical risk minimizers:
Pooling ERM:
$\widehat{f}_{\mathrm{pool}}=\underset{f\in\mathcal{G}}{\operatorname{arg\ min\ }}\frac{1}{N}\sum_{i=1}^{N}\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\ell(y_{ij},f(x_{ij})).$ (2)
Domain-informed (DI) ERM:
$\widehat{f}_{\mathrm{DG}}=\underset{f\in\mathcal{F}}{\operatorname{arg\ min\ }}\frac{1}{N}\sum_{i=1}^{N}\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\ell(y_{ij},f(x_{ij},m_{i})).$ (3)
We are interested in when DI-ERM outperforms pooling ERM. From a theoretical perspective, we compare the optimal risks associated with these two approaches.
fnum@section4 Risk and Bayes risk
To aid in understanding domain generalization, it is helpful to consider DG in relation to two other problem settings. These settings differ only in what information the classifier $f$ has access to. In all cases, the performance measure is the risk
$R(f):=\mathbb{E}_{X,Y,M,D}\left[\mathbbm{1}_{f(\cdot)\neq Y}\right],$
where the argument of $f(\cdot)$ depends on settings. In this section, we compare the risks and Bayes risks associated to these three settings. This sets the state for our theoretical results in the next section.
No Domain Information:   In this setting, the classifier only has access to the feature vector $x$ at test time, and is thus $f(x)$. As noted earlier, most empirical DG methods, especially in computer vision, have this form. The risk in this case is
$R(f)$ $=\mathbb{E}_{X,Y,M,D}\left[\mathbbm{1}_{f(X)\neq Y}\right]$
$=\mathbb{E}_{X,Y}\left[\mathbbm{1}_{f(X)\neq Y}\right]$
$=\mathbb{E}_{X}\left[\mathbb{E}_{Y|X}\left[\mathbbm{1}_{f(X)\neq Y}\right]\right],$
where, because $f$ does not depend on $D$ or $M$, these variables marginalize out. Therefore, the problem reduces to learning with respect to the marginal distribution of $(X,Y)$, which corresponds to pooling data across domains. The optimal classifier $f^{*}_{\mathrm{pool}}$ is thus the Bayes classifier for the marginal distribution of $(X,Y)$:
$f^{*}_{\mathrm{pool}}(x)=\underset{k}{\operatorname{arg\ max\ }}\mathbb{P}\left(Y=k|X=x\right).$
The corresponding Bayes risk, $R^{*}_{\mathrm{pool}}$, is the Bayes risk for the marginal distribution of $(X,Y)$:
$R^{*}_{\mathrm{pool}}:=$ $\ \mathbb{E}_{X,Y}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq Y}\right]$
$=$ $\ \mathbb{E}_{X}\left[1-\max_{k}\mathbb{P}\left(Y=k|X\right)\right].$
$R^{*}_{\mathrm{pool}}$ is the best possible performance of pooling ERM.
Partial Domain Information:   This is the setting of domain generalization. The classifier has access to not only $x$, but also the metadata $m$ that conveys partial information about the true domain $d$. A classifier in this setting is denoted $f(x,m)$. The risk is
$R(f)$ $=\mathbb{E}_{X,Y,M,D}\left[\mathbbm{1}_{f(X,M)\neq Y}\right]$
$=\mathbb{E}_{X,Y,M}\left[\mathbbm{1}_{f(X,M)\neq Y}\right]$
$=\mathbb{E}_{X,M}\left[\mathbb{E}_{Y|X,M}\left[\mathbbm{1}_{f(X,M)\neq Y}\right]\right].$
The optimal classifier $f^{*}_{\mathrm{DG}}$ is now the Bayes classifier for the distribution of $X,Y|M$,
$f^{*}_{\mathrm{DG}}(x,m)=\underset{k}{\operatorname{arg\ max\ }}\mathbb{P}\left(Y=k|X=x,M=m\right),$
and the corresponding Bayes risk is
$R^{*}_{\mathrm{DG}}:=$ $\ \mathbb{E}_{X,Y,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{DG}}(X,M)\neq Y}\right]$
$=$ $\ \mathbb{E}_{X,M}\left[1-\max_{k}\mathbb{P}\left(Y=k|X,M\right)\right].$
$R^{*}_{\mathrm{DG}}$ is the best possible performance DI-ERM. Clearly, $R^{*}_{\mathrm{pool}}\leq R^{*}_{\mathrm{DG}}$, since adding features (here, the metadata) can only decrease the Bayes risk.
Full Domain Information:   In this setting, the classifier has full knowledge of the domain $D$ at test time, and is thus denoted $f(x,d)$. In practice, full knowledge of $D$ is not available, and this setting therefore serves as a bound on the best possible performance of DG. The risk in this setting is
$R(f)$ $=\mathbb{E}_{X,Y,M,D}\left[\mathbbm{1}_{f(X,D)\neq Y}\right]$
$=\mathbb{E}_{X,Y,D}\left[\mathbbm{1}_{f(X,D)\neq Y}\right]$
$=\mathbb{E}_{X,D}\left[\mathbb{E}_{Y|X,D}\left[\mathbbm{1}_{f(X,D)\neq Y}\right]\right].$
The optimal classifier $f^{*}_{\mathrm{full}}$ is now the Bayes classifier for the distribution of $X,Y|D$,
$f^{*}_{\mathrm{full}}(x,d)=\underset{k}{\operatorname{arg\ max\ }}\mathbb{P}\left(Y=k|X=x,D=d\right),$
and the corresponding Bayes risk is:
$R^{*}_{\mathrm{full}}:=$ $\ \mathbb{E}_{X,Y,D}\left[\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq Y}\right]$
$=\ \mathbb{E}_{X,D}\left[1-\max_{k}\mathbb{P}\left(Y=k|X,D\right)\right].$
$R^{*}_{\mathrm{full}}$ serves as a lower bound for the risk in domain generalization, as described in the next section.
Remark 3
DG is very similar to multi-source domain adaptation (DA). The training data available to the learner is the same in both problems. In DA, however, the goal is to achieve $R^{*}_{\mathrm{full}}$ . Essentially, the difference between DG and DA is that DG seeks to optimize the expected risk on a new, random test domain, whereas DA seeks to optimize risk on a specific, fixed test domain. To achieve this goal, it is common in DA to make additional assumptions about how the test domain relates to the training/source domains. In DG, no such assumptions are necessary since the Bayes classifier only uses observable information.
fnum@section5 Risk Comparison
This section develops bounds that relate the three Bayes risks defined in the previous section, revealing settings where domain information is and is not beneficial. The following basic result provides a starting point.
Proposition 1 (Risk Hierarchy)
$R^{*}_{\mathrm{pool}}\geq R^{*}_{\mathrm{DG}}\geq R^{*}_{\mathrm{full}}.$
The proof is straightforward (see Section˜C.1). The first inequality is trivial, as extending a feature vector can never decrease the Bayes risk. The second inequality follows from (1).
We next describe distributional assumptions under which these inequalities become strict, and with a quantifiable gap. Under such assumptions, domain information yields improved predictive capability compared to prediction with feature vectors only. Toward that end, consider the following definition.
Definition 1 (Point-wise Margin)
Consider any random triple $(X,Y,M)$, where $Y$ is discrete. Define the point-wise margin of $Y|M=m,X=x$ as,
$\gamma(x,m):=$ $\max_{k}\mathbb{P}\left(Y=k|X=x,M=m\right)$
$-\operatorname{2nd}\max_{k}\mathbb{P}\left(Y=k|X=x,M=m\right).$
The operator $\operatorname{2nd}\max_{k}$ returns the second largest value of its argument. Thus, if the two largest values of $\mathbb{P}\left(Y=k|X=x,M=m\right)$ are the same, $\gamma(x,m)=0$. Intuitively, $\gamma(x,m)$ reflects the degree of certainty that the Bayes classifier $f^{*}_{\mathrm{DG}}(x,m)$ has about its prediction. The larger $\gamma(x,m)$, the more confident the prediction.
The next result gives upper and lower bounds on the gap between $R^{*}_{\mathrm{DG}}$ and $R^{*}_{\mathrm{pool}}$. This gap is the additional reduction in risk that results from leveraging the partial domain information $M$.
Theorem 1 (Risk Reduction w/ Domain Info)
Consider any random triple $(X,Y,M)$, where $Y$ is discrete. Then
$\mathbb{E}_{X,M}\left[\gamma(X,M)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]$
$\quad\leq R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}$
$\quad\leq\mathbb{E}_{X,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}\right].$
The proof of Theorem 1 is in Section˜C.2. The upper bound represents the probability of disagreement between the domain-informed classifier $f^{*}_{\mathrm{DG}}$ and the pooled classifier $f^{*}_{\mathrm{pool}}$. The lower bound can be interpreted as the expected cost of disagreement, where the cost is zero when the predictions agree, and equals the margin $\gamma(X,M)$ when they differ. Hence, domain information is particularly beneficial when $f^{*}_{\mathrm{DG}}$ frequently disagrees with $f^{*}_{\mathrm{pool}}$ in regions of high confidence. Theorem˜1 also implies necessary and sufficient conditions for the risk gap being zero. If $f^{*}_{\mathrm{DG}}=f^{*}_{\mathrm{pool}}$ almost surely, then $R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}=0$. If $R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}=0$, then $\gamma(X,M)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}$ must be zero almost surely, meaning $f^{*}_{\mathrm{DG}}$ and $f^{*}_{\mathrm{pool}}$ can only disagree whenever two or more classes achieve the maximum posterior probability. Figure˜1 gives more intuition.
Figure 1: Illustration of Theorem˜1. Consider binary classification with $X\in\mathbb{R}$, $Y\in\{1,2\}$, and $M\in\{1,2\}$. Then the Bayes classifiers $f^{*}_{\mathrm{pool}}(x)$, $f^{*}_{\mathrm{DG}}(x,m=1)$ and $f^{*}_{\mathrm{DG}}(x,m=2)$ can be obtained by thresholding the corresponding posteriors at $1/2$. The left figure shows a scenario where the domain-informed classifier $f^{*}_{\mathrm{DG}}$ and the pooled classifier $f^{*}_{\mathrm{pool}}$ agree everywhere, and therefore both upper and lower bound are 0. In this case, domain information $M$ is not beneficial. The right figure shows a scenario where $f^{*}_{\mathrm{DG}}$ disagrees with $f^{*}_{\mathrm{pool}}$ in certain regions, and domain information does lead to lower Bayes risk.
5.1 Decision-theoretic Gain Under Posterior Drift
Theorem 1 holds regardless of the distribution $P_{XYMD}$. By considering assumptions on this distribution, stronger conclusions may be drawn. We now examine a class of distributions where the gap $R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}$ has a more concrete lower bound. This class of distributions is motivated by applications—particularly in natural language processing—where the posterior $P_{Y\mid X,D}$ (correspondingly, $P_{Y\mid X,M}$) differs across domains due to inherent ambiguity or subjectivity. A canonical example is the sentiment or toxicity annotation task, where annotators often disagree on the same text. For instance, in the age-related sentiment analysis dataset of Díaz et al. (2018), the sentence “Old people’s appearance contains so much lived life.” received conflicting labels: 2/5 annotators seeing it as ‘very positive’, 2/5 as ‘somewhat positive’, and 1/5 as ‘very negative’. This reflects how labeling tendency varies with annotator identity. We capture this phenomenon by introducing a formal posterior drift class for domain generalization.
Definition 2 (Posterior Drift Class for DG)
$\Pi(\gamma,\epsilon):=\Big\{$ $P_{XYMD}:\forall x,m,\ \gamma(x,m)\geq\gamma,\text{ and }$
$P_{XMM^{\prime}}\Big(f^{*}_{\mathrm{DG}}\big(X,M\big)\neq f^{*}_{\mathrm{DG}}\big(X,M^{\prime}\big)\Big)\geq\epsilon\Big\},$
where $(M,M^{\prime})\mid X\sim P_{M|X}\otimes P_{M|X}$ are two independent draws.
This class of DG problems captures settings where optimal classifiers with different $M$ make conflicting predictions on a non-negligible region of the input space. The parameter $\gamma$ quantifies the point-wise confidence of the optimal predictor, and the parameter $\epsilon$ quantifies the average amount of variation in $P_{Y|X,M}$ for different $M$. With this, we have an explicit lower bound:
Proposition 2 (Gain Under Posterior Drift)
$\inf_{P_{XYMD}\in\Pi(\gamma,\epsilon)}\Big[R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}\Big]\geq\frac{\gamma\cdot\epsilon}{2}.$
The proof is given in Section˜C.3. This result establishes that, under the posterior drift class $\Pi(\gamma,\epsilon)$, the benefit of leveraging domain-specific information is guaranteed to be at least $\gamma\epsilon/2$. In other words, the risk reduction from domain-informed prediction is not only possible but provably guaranteed in this regime. This proposition can be viewed as a concrete version of the more general lower bound in Theorem˜1.
Remark 4
In contrast to pessimistic results in domain adaptation—where no method consistently outperforms vanilla ERM under posterior drift (Zhu et al., 2024; Wang et al., 2025)—our work presents an optimistic view in domain generalization: by conditioning on domain metadata $M$, we can provably do better than pooling-based prediction.
5.2 DG vs. Full Domain Knowledge
[…]
Theorem 2 ($R^{*}_{\mathrm{DG}}$ vs. $R^{*}_{\mathrm{full}}$)
Let
$\widetilde{\gamma}(x,d):=$ $\ \max_{k}\mathbb{P}\left(Y=k|X=x,D=d\right)$
$\quad-\operatorname{2nd}\max_{k}\mathbb{P}\left(Y=k|X=x,D=d\right).$
Then
$\mathbb{E}_{X,D,M}\left[\widetilde{\gamma}(X,D)\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]$
$\quad\leq R^{*}_{\mathrm{DG}}-R^{*}_{\mathrm{full}}$
$\quad\leq\mathbb{E}_{X,D,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq f^{*}_{\mathrm{DG}}(X,M)}\right].$
The result parallels Theorem˜1: if domain information is low quality—i.e., $f^{*}_{\mathrm{full}}$ often disagrees with $f^{*}_{\mathrm{DG}}$ with high-confidence—then $R^{*}_{\mathrm{DG}}$ can be much larger than $R^{*}_{\mathrm{full}}$. The necessary and sufficient conditions for this gap to vanish are analogous to those in Theorem˜1.
5.3 Gain Under Covariate Shift?
Covariate shift refers to a setting where, as the domain information $M$ varies, $P_{X|M}$ changes, but $P_{Y|X,M}$ does not (Quiñonero-Candela et al., 2008). More generally, we may extend the meaning of covariate shift to be any DG problem where the Bayes classifier $f^{*}_{\mathrm{DG}}$ does not depend on $M$. In such a scenario, access to domain-specific information cannot reduce the Bayes risk.
Corollary 1
Under covariate shift, $R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DG}}$.
Nevertheless, empirical studies have reported performance gains from incorporating domain information even under covariate shift (Dubey et al., 2021; Bui et al., 2021). Our next result shows that such gains do not reflect any decision-theoretic benefit, but instead arise from the restrictions of the chosen function class.
Assumption 1 (DI & Pooling Function Classes)
Let $\{\mathcal{F}_{k}\}_{k=1}^{\infty}\subset\{\mathcal{X}\times\mathcal{M}\to\mathcal{Y}\}$ and $\{\mathcal{G}_{k}\}_{k=1}^{\infty}\subset\{\mathcal{X}\to\mathcal{Y}\}$ be sequences of function classes related as follows:
  1. 1.
    For every $g\in\mathcal{G}_{k}$, there exists $f\in\mathcal{F}_{k}$ and some fixed $m_{0}\in\mathcal{M}$ such that
    $g(x)=f(x,m_{0}).$
  2. 2.
    The pooling classes $\{\mathcal{G}_{k}\}$ satisfy the universal approximation property: for any distribution $P_{XY}$,
    $\inf_{g\in\mathcal{G}_{k}}R(g)\;\to\;R^{*}_{\mathrm{pool}}\quad\text{as }k\to\infty.$
The first condition is satisfied by common architectures (e.g., feedforward ReLU networks) and $\mathcal{G}$ can be induced from $\mathcal{F}$ by setting the weights from the $m$ inputs to zero. What’s more, $\mathcal{G}$ satisfies the universal approximation property as the network size grows.
Proposition 3
For function classes $\mathcal{F}_{k}$ and $\mathcal{G}_{k}$ satisfying Assumption 1, denote
$R^{*}_{\mathrm{pool},\mathcal{G}_{k}}:=\inf_{g\in\mathcal{G}_{k}}R(g)\ \text{ and }\ R^{*}_{\mathrm{DG},\mathcal{F}_{k}}:=\inf_{f\in\mathcal{F}_{k}}R(f).$
Under covariate shift,
$R^{*}_{\mathrm{pool},\mathcal{G}_{k}}\geq R^{*}_{\mathrm{DG},\mathcal{F}_{k}}\quad\forall k,$ (4)
and
$\lim_{k\to\infty}R_{\mathrm{pool},\mathcal{G}_{k}}=\lim_{k\to\infty}R_{\mathrm{DG},\mathcal{F}_{k}}.$ (5)
This shows that with a restricted function class $\mathcal{G}_{k}$, leveraging domain-information may lead to a benefit under covariate shift. Furthermore, strict inequality in Equation˜4 is possible, as illustrated with a simple example in Appendix˜B. However, these gains vanish asymptotically as the function classes become more expressive. This phenomenon is also observed empirically in Section˜6.
fnum@section6 Experiments
We evaluate the effectiveness of domain-informed ERM (DI-ERM) in three experimental settings. Our primary focus is on the comparison between DI-ERM and pooling ERM, which highlights the benefit of incorporating domain metadata. Additional results—including linear probing, benchmarks against alternative and state-of-the-art methods, and complete experimental details—are provided in Appendix˜D.
6.1 Annotator Disagreement
In many NLP tasks, annotators exhibit subjective preferences, leading to disagreement on the label $y$ for the same input xx—a form of posterior drift discussed in Section˜5.1. To study this phenomenon, we use the dataset of Díaz et al. (2018), which re-annotates a subset of the Sentiment140 dataset (Go et al., 2009) for training and provides a test set drawn from age-related blog posts. The training set comprises 59,235 sentences labeled by 1,481 annotators; the test set includes 1,419 sentences labeled by 878 annotators. Each sentence is annotated by 4–5 individuals, and the labels exhibit high disagreement (about 40 $\%$). In this setting, the input $x$ is a sentence, the label $y\in\{1,2,3,4,5\}$ denotes sentiment on a five-point scale, the domain $d$ corresponds to the annotator, and the domain information $m$ consists of demographic metadata (e.g., age, upbringing region).
To encode domain information $M$, we concatenate it with the sentence $x$ in a text-prompt format, as illustrated in Figure˜3. We finetune the BERT base model (Devlin et al., 2019). Table˜2 reports the results. Since prior DG methods have not been evaluated on this dataset, we implemented and benchmarked them under the same experimental setup. DI-ERM substantially outperforms pooling ERM, demonstrating that leveraging annotator metadata can dramatically improve predictive accuracy.
Table 2: Test accuracy on the sentiment disagreement dataset. Previous DG methods had not been benchmarked on this dataset; we re-implemented them for comparison under our setup. Incorporating annotator profiles ($M$) through DI-ERM yields a dramatic improvement over pooling ERM, reflecting the importance of modeling annotator-specific posterior drift. In particular, DI-ERM nearly doubles accuracy compared to pooling ERM and surpasses the previous state-of-the-art by Deng et al. (2023).
Algorithm Model Test Avg Acc
Pooling ERM BERT 49.1 $\pm$ 0.4
IRM BERT 48.1 $\pm$ 0.7
GroupDRO BERT 49.1 $\pm$ 0.1
CORAL BERT 48.4 $\pm$ 0.2
AnnEmb (SOTA) BERT 64.6 $\pm$ 0.8
DI-ERM (ours) BERT 90.5 $\pm$ 0.2
6.2 Reviewer-specific Analysis
We next examine the WILDS-Amazon Reviews dataset (Koh et al., 2021), which captures distributional shifts across reviewers. Here, the input $x$ is a product review, $y\in\{1,\dots,5\}$ is the star rating, $d$ denotes the reviewer identity, and $m$ consists of all (unlabeled) reviews written by that reviewer. Once again, because of differences among reviewers, posterior drift is exhibited.
The central hypothesis is that a reviewer’s writing style $M=P_{X|D}$ provides a useful signal for predicting their rating behavior $P_{Y|X,D}$. The training set contains 245,502 reviews from 1,252 reviewers, while the test set consists of 100,050 reviews from 1,334 unseen reviewers.
To incorporate reviewer context $M$, we randomly sample 20 additional reviews written by the same reviewer and concatenate them with the current review in a prompt format, shown in Figure˜4. We finetune the nomic-embed-text-v1.5 model (Nussbaum et al., 2025), which supports a longer context window than BERT so that domain information $M$ can fit in. As summarized in Table˜3, DI-ERM outperforms pooling ERM. Beyond higher average accuracy, DI-ERM also boosts the 10th-percentile accuracy across reviewers—a key robustness metric used on the official leaderboard.
[…]
Algorithm Model Test Avg Acc Test 10% Acc
Pooling ERM DistilBERT 72.0 $\pm$ 0.1 54.2 $\pm$ 0.8
GroupDRO DistilBERT 70.0 $\pm$ 0.5 53.3 $\pm$ 0.8
CORAL DistilBERT 71.1 $\pm$ 0.3 52.9 $\pm$ 0.8
IRM DistilBERT 70.3 $\pm$ 0.6 52.4 $\pm$ 0.8
LISA (SOTA) DistilBERT 70.7 $\pm$ 0.3 54.7 $\pm$ 0.0
Pooling ERM nomic-embed 72.2 $\pm$ 1.1 55.1 $\pm$ 0.8
DI-ERM (ours) nomic-embed 73.4 $\pm$ 0.1 56.4 $\pm$ 0.8
6.3 Image Classification Across Styles
We next evaluate our method on the PACS dataset (Li et al., 2017), which contains images from four distinct visual styles (aka, domain $d$): Photo (P), Art Painting (A), Cartoon (C), and Sketch (S). Each image $x$ belongs to one of seven categories: dog, elephant, giraffe, guitar, horse, house, or person. Domain information $m$ is encoded by a short text description $m$, such as “a photo” or “a pencil sketch” (see Figure˜5).
This vision task satisfies covariate shift, since a single classifier should accurately classify all images across domains. Thus, in line with Section˜5.3, we expect any gains to be due to using a restricted function class. To implement DI-ERM, we use pretrained image foundation models (e.g., CLIP (Radford et al., 2021), DINOv2 (Oquab et al., 2023), DINOv3 (Siméoni et al., 2025)) to extract visual features from $x$, and encode the domain description $m$ using a pretrained language model (DistilBERT) following the prompt in Figure˜5. The resulting image and text embeddings are concatenated into a joint representation for classification.
We follow the standard PACS evaluation protocol: training on three domains and testing on the held-out fourth domain, repeated across all domain splits. All encoders are frozen, and linear classifiers are trained on top of the fixed representations.
As shown in Section˜6.3, DI-ERM improves over pooling ERM in most settings. The gains are most pronounced for mid-sized models, while the benefit diminishes for larger foundation models. This pattern aligns with the discussion in Section˜5.3: under covariate shift, the benefit of DI-ERM decreases as model size grows.
Table 4: Domain generalization results on PACS using models from the CLIP, DINOv2 and DINOv3 families. DI-ERM achieves improved accuracy over pooling ERM in most configurations, moreso for mid-sized models than for large models, a phenomenon consistent with our theory in Section˜5.3. When using large models, both ERM and DI-ERM approach SOTA performance by Cho et al. (2023).
Algorithm Model Test Avg Acc
PromptStyler (SOTA) CLIP: ViT-large 98.6 $\pm$ 0.0
Pooling ERM (linear) CLIP: ViT-small 95.0 $\pm$ 0.0
DI-ERM (linear) 95.4 $\pm$ 0.0
gray!80blackPooling ERM (linear) CLIP: ViT-large 98.6 $\pm$ 0.0
DI-ERM (linear) 98.6 $\pm$ 0.0
Pooling ERM (linear) DINOv2: ViT-small 87.6 $\pm$ 0.0
DI-ERM (linear) 89.0 $\pm$ 0.0
gray!80blackPooling ERM (linear) DINOv2: ViT-large 96.1 $\pm$ 0.0
DI-ERM (linear) 96.3 $\pm$ 0.0
Pooling ERM (linear) DINOv3: ViT-small 91.7 $\pm$ 0.0
DI-ERM (linear) 93.3 $\pm$ 0.0
gray!80blackPooling ERM (linear) DINOv3: ViT-large 95.4 $\pm$ 0.0
DI-ERM (linear) 96.2 $\pm$ 0.0
fnum@section7 Conclusions
This work presents a rigorous theory of domain generalization, precisely characterizing when and why leveraging domain information at test time is beneficial. Empirically, we demonstrate that domain-informed ERM (DI-ERM) outperforms pooled ERM across three representative scenarios in language and vision tasks. A key requirement of DI-ERM is access to domain metadata; when such metadata is unavailable, unlabeled data from the target domain offers a natural alternative.
Future work might explore alternative ways of encoding domain information, and a broader range of DG benchmarks. Additionally, it would interesting to explore the phenomena discussed in the context of other performance measures for DG, such as the worst-case test error examined in (Dwork et al., 2025).
Acknowledgements
We thank Gilles Blanchard for helpful discussions and insights.
Supplementary Materials
Appendix A Additional literature review
ERM is hard to beat.
Empirically, Gulrajani and Lopez-Paz (2021) first emphasized that a well-tuned empirical risk minimization (ERM) baseline outperforms many domain generalization (DG) methods on vision benchmarks. Similar patterns were later observed on the WILDS benchmark (Koh et al., 2021), and again in the context of federated domain generalization by Bai et al. (2024).
On the theoretical front, Rosenfeld et al. (2021) and Gouk et al. (2024) studied function classes of the form $f:\mathcal{X}\to\mathcal{Y}$ and concluded that, under common assumptions, ERM cannot be fundamentally outperformed (e.g., in terms of minimax risk). The recent work by Dwork et al. (2025) analyzed the optimality of the “min-max” ERM with another performance measure of interest.
The use of unlabeled data.
While most DG methods restrict themselves to using only the input $x$ at inference time, some methods explore the use of unlabeled test-domain data. Several DG methods attempt to exploit unlabeled test data to improve generalization (Blanchard et al., 2011; Muandet et al., 2013; Zhang et al., 2021). A closely related setting is unsupervised domain adaptation (UDA), where unlabeled test data are used to adapt models to the test domain. Unlike DG, UDA assumes access to target-domain data at training time and typically requires learning a separate model per test domain (Sun and Saenko, 2016; Ganin et al., 2016).
Although promising in principle, the practical benefits of using unlabeled data remain mixed. A large-scale study by Sagawa et al. (2022) evaluating methods across ten diverse datasets found that incorporating unlabeled data frequently failed to improve upon strong ERM baselines. These findings reinforce the need for a more precise understanding of when and how unlabeled data can contribute to domain generalization.
Our framework addresses this gap by casting unlabeled data as a special case of auxiliary domain information, and by providing conditions under which such information is expected to improve generalization performance.
Appendix B A Simple Example of Covariate Shift
The example below show that even if there is no decision-theoretic gain of DI-ERM under covariate shift (Corollary˜1), it may still have practical advantage when considering learning over a restricted function class.
Figure 2: Illustration of Example 1, where $R^{*}_{\mathrm{pool},\,\mathcal{G}}>R^{*}_{\mathrm{DG},\,\mathcal{F}}$.
Example 1 (Covariate shift without posterior drift)
Let $P_{XYM}$ be
$M\sim\mathrm{Bernoulli}(p),\text{ where }p>1/2,$
$\begin{cases}M=1:&X\!\sim\!\mathrm{Unif}[0,2],\;Y=\operatorname{sign}(X-1)\\ M=2:&X\!\sim\!\mathrm{Unif}[4,6],\;Y=\operatorname{sign}(X-5).\end{cases}$
Because the supports are disjoint, the pooling and DG Bayes classifier are the same, to be specific
$f^{*}_{\mathrm{pool}}(x)$ $=\begin{cases}\operatorname{sign}(x-1),&x\in[0,2]\\ \operatorname{sign}(x-5),&x\in[4,6]\end{cases},$
$f^{*}_{\mathrm{DG}}(x,m)$ $=\operatorname{sign}(x-4m+3)\quad\Longrightarrow f^{*}_{\mathrm{pool}}=f^{*}_{\mathrm{DG}}$
therefore $R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DG}}=0$. The model classes are linear classifiers
$\mathcal{F}$ $=\bigl\{f(x,m)=\operatorname{sign}(w^{\top}x+v^{\top}m+b)\bigr\},$
$\mathcal{G}$ $=\bigl\{f(x)=\operatorname{sign}(w^{\top}x+b)\bigr\}.$
$\mathcal{F}$ can realize $f^{*}_{\mathrm{DG}}$ with a bias term that depends on $m$, giving $R^{*}_{\mathrm{DG},\,\mathcal{F}}=0$. However, a predictor in $\mathcal{G}$ can only choose a single threshold, and the optimal one is
$f^{*}_{\mathrm{pool},\,\mathcal{G}}(x)=\operatorname{sign}(x-1),\quad R^{*}_{\mathrm{pool},\,\mathcal{G}}=\frac{\min\{p,1-p\}}{2}.$
Therefore, $R^{*}_{\mathrm{pool},\,\mathcal{G}}>R^{*}_{\mathrm{DG},\,\mathcal{F}}$, even though $R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DG}}$.
This toy construction mirrors image classification task across different styles: each style (domain) has a separate support, so the Bayes classifier is the same with or without $m$, yet $m$ still helps within a restricted model class. This is experimentally illustrated in Section˜6 .
Appendix C Proofs
C.1 Proof of Proposition˜1
Proposition (Risk Hierarchy)  $R^{*}_{\mathrm{pool}}\geq R^{*}_{\mathrm{DG}}\geq R^{*}_{\mathrm{full}}.$
Proof. Although this result follows from Theorems 1 and 2, it can be proved more directly as follows:
$R^{*}_{\mathrm{pool}}$ $=\inf_{f:\mathcal{X}\to\mathcal{Y}}\mathbb{E}_{X,Y,M,D}\left[\mathbbm{1}_{f(X)\neq Y}\right]$
$\geq\inf_{f:\mathcal{X}\times\mathcal{M}\to\mathcal{Y}}\mathbb{E}_{X,Y,M,D}\left[\mathbbm{1}_{f(X,M)\neq Y}\right]=R^{*}_{\mathrm{DG}}$
$\geq\inf_{f:\mathcal{X}\times\mathcal{M}\times\mathcal{D}\to\mathcal{Y}}\mathbb{E}_{X,Y,M,D}\left[\mathbbm{1}_{f(X,M,D)\neq Y}\right]$
$=\inf_{f:\mathcal{X}\times\mathcal{D}\to\mathcal{Y}}\mathbb{E}_{X,Y,M,D}\left[\mathbbm{1}_{f(X,D)\neq Y}\right]=R^{*}_{\mathrm{full}}.$
To see the final step, note that the $f(x,m,d)$ achieving the minimum in the next-to-last line is $\mathop{\mathrm{argmax}}_{k}{\mathbb{P}}(Y=k\mid X=x,M=m,D=d)$. By Assumption (1), this is equal to $\mathop{\mathrm{argmax}}_{k}{\mathbb{P}}(Y=k\mid X=x,D=d)$, which achieves the minimum in the last line.
C.2 Proof of Theorem˜1
Theorem (Risk Reduction from Domain Information)  Consider any random triple $(X,Y,M)$, where $Y$ is discrete. Then
$\mathbb{E}_{X,M}\left[\gamma(X,M)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]\leq R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}\leq\mathbb{E}_{X,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]$
and in particular,
$f^{*}_{\mathrm{pool}}(X)=f^{*}_{\mathrm{DG}}(X,M)\quad\text{ almost surely w.r.t. }P_{XM}\implies R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DG}}.$
Proof. The gap in the two risks can be expressed as
$R^{*}_{\mathrm{pool}}$ $-R^{*}_{\mathrm{DG}}$
$=$ $\ \mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(X,M)|X,M\right)\right]\right]-\mathbb{E}_{X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(X)|X\right)\right]$
$=$ $\ \mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(X,M)|X,M\right)\right]\right]-\mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(X)|X,M\right)\right]\right]$
$=$ $\ \mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(X,M)|X,M\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(X)|X,M\right)\right]\right]$
Notice that for any $x,m$, if $f^{*}_{\mathrm{DG}}(x,m)=f^{*}_{\mathrm{pool}}(x)$, then the pointwise difference of the conditional probabilities inside the expectation above must be zero.
Whereas if they disagree, then it must hold that
$\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(X,M)|X,M\right)$ $=\max_{k}P(Y=k|X,M),$
and
$\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(X)|X,M\right)$ $\leq\operatorname{2nd}\max_{k}P(Y=k|X,M).$
It thus follows that.
$\gamma(x,m)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DG}}(x,m)}$ $\leq\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(x,m)|X=x,M=m\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(x)|X=x,M=m\right)$
$\leq\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DG}}(x,m)}.$
The inequalities in the theorem statement now follow.
From the lower and upper bound above, we can directly get the sufficient condition that
$f^{*}_{\mathrm{pool}}(X)=f^{*}_{\mathrm{DG}}(X,M)\quad\text{ almost surely w.r.t. }P_{XM}\implies R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DG}}.$
C.3 Proof of Proposition˜2
Proposition
$\inf_{P_{XYMD}\in\Pi(\gamma,\epsilon)}\Big[R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}\Big]\geq\frac{\gamma\cdot\epsilon}{2}$
Proof. From the lower bound in Theorem˜1, we have
[…]
$\geq\gamma\ \mathbb{E}_{X,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]$ $\because\text{margin assumption in }\Pi(\gamma,\epsilon)$
$=\gamma\ \mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]\right]$
Now we will show that
$\forall x,\quad\mathbb{E}_{M|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DG}}(x,M)}\right]\geq\frac{1}{2}\ \mathbb{E}_{M,M^{\prime}|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{DG}}(x,M)\neq f^{*}_{\mathrm{DG}}(x,M^{\prime})}\right],$
where
$M,\ M^{\prime}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}P_{M|X=x}.$
Let’s examine the two terms. Fix $x$, denote
$\pi_{k}(x)=\mathbb{P}\left(f^{*}_{\mathrm{DG}}(x,M)=k|X=x\right),$
and note that the randomness comes from $M$.
Then for any $x$,
$\mathbb{E}_{M,M^{\prime}|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{DG}}(x,M)\neq f^{*}_{\mathrm{DG}}(x,M^{\prime})}\right]$ $=\mathbb{P}\left(f^{*}_{\mathrm{DG}}(x,M)\neq f^{*}_{\mathrm{DG}}(x,M^{\prime})|X=x\right)$
$=\sum_{k}\mathbb{P}\left(f^{*}_{\mathrm{DG}}(x,M)=k,f^{*}_{\mathrm{DG}}(x,M^{\prime})\neq k|X=x\right)$
$=\sum_{k}\pi_{k}(x)\left(1-\pi_{k}(x)\right)$
$=1-\sum_{k}\pi_{k}(x)^{2}.$
Now assume $f^{*}_{\mathrm{pool}}(x)=k_{0}$. Then
$\mathbb{E}_{M|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DG}}(x,M)}\right]$ $=\mathbb{E}_{M|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{DG}}(x,M)\neq k_{0}}\right]$
$=1-\pi_{k_{0}}(x)$
Notice that
$1-\sum_{k}\pi_{k}(x)^{2}$ $\leq 1-\pi_{k_{0}}^{2}$ =” $\textbf{``$=$'' when }\pi_{k_{0}}=1$
$=(1+\pi_{k_{0}})(1-\pi_{k_{0}})$
$\leq 2(1-\pi_{k_{0}})$ =” $\textbf{``$=$'' when }\pi_{k_{0}}=1.$
Then
$\mathbb{E}_{M|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DG}}(x,M)}\right]\geq\frac{1}{2}\ \mathbb{E}_{M,M^{\prime}|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{DG}}(x,M)\neq f^{*}_{\mathrm{DG}}(x,M^{\prime})}\right].$
Integrating over $x$, we have
$\mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]\right]$ $\geq\frac{1}{2}\mathbb{E}_{X}\left[\mathbb{E}_{M,M^{\prime}|X}\left[\mathbbm{1}_{f^{*}_{\mathrm{DG}}(X,M)\neq f^{*}_{\mathrm{DG}}(X,M^{\prime})}\right]\right]$
$=\frac{1}{2}P_{X,M,M^{\prime}}\Big(f^{*}_{\mathrm{DG}}\big(X,M\big)\neq f^{*}_{\mathrm{DG}}\big(X,M^{\prime}\big)\Big)$
$\geq\frac{1}{2}\epsilon.\qquad\because\text{by definition of }\Pi(\gamma,\epsilon)$
C.4 Proof of Theorem˜2
Theorem  Let
$\widetilde{\gamma}(x,d):=\max_{k}\mathbb{P}\left(Y=k|X=x,D=d\right)-\operatorname{2nd}\max_{k}\mathbb{P}\left(Y=k|X=x,D=d\right)$
Then
$\mathbb{E}_{X,D,M}\left[\gamma(X,D)\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]\leq R^{*}_{\mathrm{DG}}-R^{*}_{\mathrm{full}}\leq\mathbb{E}_{X,D,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{full}}(X,D)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]$
and in particular,
$f^{*}_{\mathrm{full}}(X,D)=f^{*}_{\mathrm{DG}}(X,M)\quad\text{ almost surely w.r.t. }P_{XMD}\implies R^{*}_{\mathrm{full}}=R^{*}_{\mathrm{DG}}.$
Proof.
$R^{*}_{\mathrm{DG}}-R^{*}_{\mathrm{full}}$ $=\mathbb{E}_{X,Y,D,M}\left[\mathbbm{1}_{Y\neq f^{*}_{\mathrm{full}}(X,D)}\right]-\mathbb{E}_{X,Y,D,M}\left[\mathbbm{1}_{Y\neq f^{*}_{\mathrm{DG}}(X,M)}\right]$
$=\mathbb{E}_{X,D,M}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(X,M)\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{full}}(X,D)\right)|X,D,M\right]$
By Assumption (1):
$Y|X,D,M=Y|X,D.$
Then for every $x$, $d$ and $m$,
$\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(x,m)|X=x,D=d,M=m\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{full}}(x,d)|X=x,D=d,M=m\right)$
$=\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(x,m)|X=x,D=d\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{full}}(x,d)|X=x,D=d\right)$
$\geq\widetilde{\gamma}(x,d)\mathbbm{1}_{f^{*}(x,m)\neq f^{*}(x,d)}.$
In the other direction,
$\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(x,m)|X=x,D=d,M=m\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{full}}(x,d)|X=x,D=d,M=m\right)$
$\leq\mathbbm{1}_{f^{*}(x,m)\neq f^{*}(x,d)}.$
Integrating over $x,d,m$, we get the lower and upper bounds.
From the lower and upper bound, we can directly get the sufficient condition
$f^{*}_{\mathrm{full}}(x,d)=f^{*}_{\mathrm{DG}}(x,m)\quad\text{ almost surely w.r.t. }P_{XMD}.\implies R^{*}_{\mathrm{full}}=R^{*}_{\mathrm{DG}}$
C.5 Proof of Proposition˜3
Proposition  For function classes $\{\mathcal{F}_{k}\}$ and $\{\mathcal{G}_{k}\}$ satisfying Assumption 1. Under covariate shift,
$R_{\mathrm{pool},\mathcal{G}_{k}}:=\inf_{g\in\mathcal{G}_{k}}R(g)\geq R_{\mathrm{DG},\mathcal{F}_{k}}:=\inf_{f\in\mathcal{F}_{k}}R(f),\quad\forall k,$ (6)
and
$\lim_{k\to\infty}R_{\mathrm{pool},\mathcal{G}_{k}}=\lim_{k\to\infty}R_{\mathrm{DG},\mathcal{F}_{k}}$ (7)
Proof. The first line follows from the first assumption of of Assumption 1, where every function $g\in\mathcal{G}_{k}$ is also realizable in $\mathcal{F}_{k}$.
The second line follows from covariate shift assumption
$R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DG}}$
the universal approximation property of function class $\mathcal{G}_{k}$
$\lim_{k\to\infty}R_{\mathrm{pool},\mathcal{G}_{k}}=R^{*}_{\mathrm{pool}},$
and the fact that
$R^{*}_{\mathrm{DG}}\leq R_{\mathrm{DG},\mathcal{F}_{k}}\leq R_{\mathrm{pool},\mathcal{G}_{k}}.$
Appendix D Experimental details
This section provides additional details on our experimental setup, models, and performance comparisons. Unless otherwise specified, all models used for fine-tuning are implemented using publicly available checkpoints (e.g., via Huggingface, Pytorch, or official Github repo). For linear probing experiments, we extract feature representations using pre-trained transformers and train downstream classifiers with scikit-learn, using either logistic regression or multilayer perceptrons (MLPs).
The following subsections follows the same structure as Section˜6, while providing additional details and full tables.
D.1 Annotator disagreement
Fine-Tuning.
We fine-tune the bert-base-uncased model and benchmark DI-ERM against other domain generalization methods. For DI-ERM, we concatenate the sentence $x$ with the annotator profile $m$ using the text prompt shown in Figure˜3.
All experiments are done on NVIDIA A100 or NVIDIA A40 GPU. We finetune the model with logistic loss for 10 epochs with learning rate $5\cdot 10^{-5}$. Table˜5 reports the results over three trials. Our models consistently outperform prior work, with the best configuration achieving over $90\%$ test accuracy—substantially higher than the previous state-of-the-art reported by Deng et al. (2023).
Instruction: Read the following sentence and the annotator’s demographic profile and determine how positive or negative the annotator judged the sentence on a 1--5 scale (1 = Very negative, 5 = Very positive). Sentence: [sentence goes here] Annotator profile: Age {age}, Race {race}, Hispanic/Latino {hisp}, grew up in {grew}, currently lives in {curr}, region {region}, income {income}, education {education}, employment {employment}, living situation {living}, politics {politics}, gender {gender}. Answer:
Figure 3: Text prompt that encodes annotator profile.
Table 5: Test accuracy on the sentiment disagreement dataset (fine-tuning BERT). DI-ERM (ours) achieves the best performance.
Algorithm Model Test Avg Acc
ERM BERT 49.1 $\pm$ 0.4
IRM BERT 48.1 $\pm$ 0.7
GroupDRO BERT 49.1 $\pm$ 0.1
CORAL BERT 48.4 $\pm$ 0.2
AnnEmb (SOTA) BERT 64.6 $\pm$ 0.8
DI-ERM (ours, fine-tune) BERT 90.5 $\pm$ 0.2
Linear/MLP-probing.
We also evaluate in a frozen-feature setting, where the language model is fixed and a lightweight classifier is trained on top. Here, $x$ is encoded with a pretrained sentiment model (e.g., [CLS] embedding of DistilBERT checkpoint that fine-tuned on SST-2 dataset), while $m$ is encoded with the DistilBERT base model. The embeddings are concatenated and passed to either a linear or shallow MLP classifier. The classifiers are trained in scikit-learn.
Section˜D.1 presents the results. DI-ERM consistently outperforms pooling ERM across different feature extractors.
Table 6: Test accuracy on the sentiment disagreement dataset (frozen feature extractor). DI-ERM consistently outperforms pooling ERM, and in some settings surpasses the prior state-of-the-art of Deng et al. (2023). We highlight the best performance reported by Deng et al. (2023) (69.77) and our highest score (83.41). $\dagger$: Checkpoints used in Deng et al. (2023) were not publicly specified.
Algorithm Model Test Avg Acc
AnnEmb (Deng et al., 2023) BERT 64.61
RoBERTa 60.30
DeBERTa 69.77
Pooling ERM (linear) distilbert-base-uncased-finetuned-sst-2-english 45.85
DI-ERM (linear) distilbert-base-uncased-finetuned-sst-2-english 46.42
gray!80blackPooling ERM (MLP) distilbert-base-uncased-finetuned-sst-2-english 55.07
DI-ERM (MLP) distilbert-base-uncased-finetuned-sst-2-english 78.45
Pooling ERM (linear) bert-base-multilingual-uncased-sentiment 43.06
DI-ERM (linear) bert-base-multilingual-uncased-sentiment 43.94
gray!80blackPooling ERM (MLP) bert-base-multilingual-uncased-sentiment 53.90
DI-ERM (MLP) bert-base-multilingual-uncased-sentiment 83.41
D.2 Reviewer-specific analysis
Fine-Tuning.
We fine-tune the bert-base-uncased model and benchmark DI-ERM against other domain generalization methods. For DI-ERM, we concatenate each review $x$ with reviewer context $m$, represented by 20 randomly selected reviews from the same reviewer, using the text prompt in Figure˜4.
We choose nomic-embed-text-v1.5 (Nussbaum et al., 2025), which supports a context window of up to 8012 tokens (we choose 2048 in our experiments), in order to handle the long reviwer context $m$.
All experiments are done on NVIDIA A100 or NVIDIA A40 GPU. We finetune the model with logistic loss for 2 epochs with learning rate $10^{-5}$. Table˜7 reports the results over three trials. DI-ERM achieves the best performance, outperforming previously reported methods on the WILDS leaderboard (https://wilds.stanford.edu/).
Instruction: Classify the current review based on this reviewer’s sentiment patterns. Current Review: [current review goes here] Reviewer’s Historical Reviews: Review 1: [review_1]   |   Review 2: [review_2]   |   …
Figure 4: Text prompt that encodes reviewer writing style
Table 7: Reviewer-specific sentiment analysis. DI-ERM (ours) achieves the highest accuracy, outperforming prior state-of-the-art by Yao et al. (2022).
Algorithm Model Test Avg Acc Test 10% Acc
ERM DistilBERT 72.0 $\pm$ 0.1 54.2 $\pm$ 0.8
GroupDRO DistilBERT 70.0 $\pm$ 0.5 53.3 $\pm$ 0.8
CORAL DistilBERT 71.1 $\pm$ 0.3 52.9 $\pm$ 0.8
IRM DistilBERT 70.3 $\pm$ 0.6 52.4 $\pm$ 0.8
LISA (SOTA) DistilBERT 70.7 $\pm$ 0.3 54.7 $\pm$ 0.0
ERM nomic-embed-text 72.2 $\pm$ 1.1 55.1 $\pm$ 0.8
DI-ERM (ours) nomic-embed-text 73.4 $\pm$ 0.1 56.4 $\pm$ 0.8
Linear/MLP-probing.
We also evaluate in a frozen-feature setting, where the language model is fixed and only a lightweight classifier is trained. Each review $x$ is represented by its [CLS] embedding from a pretrained sentiment model (e.g., DistilBERT fine-tuned on SST-2). For reviewer context $m$, we average the [CLS] embeddings of all reviews written by that reviewer. The concatenated review and reviewer embeddings are then passed to a linear or a shallow MLP classifier implemented in scikit-learn.
Domain2Vec.
Inspired by Zaheer et al. (2017); Deshmukh et al. (2018), we implement a Domain2Vec-style module to encode reviewer-specific domain information. Given a set of reviews $\{x_{1},x_{2},\dots,x_{n}\}\sim P_{X|D=d}$ written by reviewer $d$, we learn a mapping
$f(\{x_{1},x_{2},\dots,x_{n}\})=\rho\left(\frac{1}{n}\sum_{i=1}^{n}\phi(x_{i})\right),$
where $\phi$ and $\rho$ are MLPs that map individual feature representations (extracted from pretrained model) to a latent space and then transform the aggregated feature, respectively. The resulting vector is concatenated with the review representation $x$ to predict its sentiment label $y$.
Section˜D.2 shows the result.
Table 8: Sentiment classification on Amazon-WILDS with reviewer-specific signals. “Domain2Vec” denotes reviewer encoding based on a learned mean embedding. DI-ERM variants consistently outperform pooling ERM baselines.
Algorithm Model Test Avg Acc Test 10% Acc
Pooling ERM (linear) distilbert-base-uncased-finetuned-sst-2-english 67.42 48.00
DI-ERM (linear) distilbert-base-uncased-finetuned-sst-2-english 68.21 48.00
gray!80blackPooling ERM (MLP) distilbert-base-uncased-finetuned-sst-2-english 67.59 48.00
DI-ERM (MLP) distilbert-base-uncased-finetuned-sst-2-english 68.28 49.33
gray!80blackDI-ERM (Domain2Vec) distilbert-base-uncased-finetuned-sst-2-english 68.21 48.00
Pooling ERM (linear) bert-base-multilingual-uncased-sentiment 72.14 53.33
DI-ERM (linear) bert-base-multilingual-uncased-sentiment 73.22 54.67
gray!80blackPooling ERM (MLP) bert-base-multilingual-uncased-sentiment 73.01 53.33
DI-ERM (MLP) bert-base-multilingual-uncased-sentiment 73.18 55.07
gray!80blackDI-ERM (Domain2Vec) bert-base-multilingual-uncased-sentiment 73.19 54.67
D.3 Image classification across styles
We evaluate our approach on the PACS benchmark, which contains four visual styles: Photo (P), Art Painting (A), Cartoon (C), and Sketch (S). To assess robustness to style variation, we test a diverse set of models from the CLIP, DINOv2 and DINOv3 families.
For all the experiment we use the text prompt in Figure˜5 as input to DistillBERT.
Domain "photo", text prompt: "a photo" Domain "art painting", text prompt: "an oil painting" Domain "cartoon", text prompt: "a colorful cartoon" Domain "sketch", text prompt: "a pencil sketch"
Figure 5: Example of style-specific text prompts used as domain descriptions.
We use different models to extract the visual features. Then train a logistic regression on top of these features using Scikit-learn with lbfgs solver, the results are deterministic. Section˜D.3 summarizes the results.
Across most domain shifts, our proposed DI-ERM method consistently outperforms standard pooling ERM, highlighting the advantage of incorporating domain-specific information into the representation.
Notably, we observe that the performance gains from DI-ERM tend to diminish as model capacity increases. For the largest models (e.g., CLIP ViT-L/14 and DINOv2 ViT-L/14), the improvement is marginal or saturates. This trend is also observed by various empirical works, e.g. Cho et al. (2023).
Table 9: Domain generalization results on PACS using models from the CLIP, DINOv2 and DINOv3 families. DI-ERM achieves improved accuracy over pooling ERM in most configurations, particularly for mid-sized models.
Model Algorithm PAC $\to$ S ACS $\to$ P CSP $\to$ A SPA $\to$ C Test Avg Acc
CLIP: vitb32 Pooling ERM (linear) 86.97 99.58 95.90 97.48 94.98
DI-ERM (linear) 88.06 99.64 96.29 97.48 95.37
gray!80blackCLIP: vitb16 Pooling ERM (linear) 90.89 99.70 97.51 98.76 96.70
DI-ERM (linear) 91.09 99.70 97.61 98.76 96.79
gray!80blackCLIP: vitl14 Pooling ERM (linear) 95.42 99.94 99.22 99.79 98.59
DI-ERM (linear) 95.32 99.94 99.32 99.79 98.59
DINOv2: vits14 Pooling ERM (linear) 79.82 85.81 93.55 91.34 87.63
DI-ERM (linear) 80.45 90.00 94.09 91.60 89.04
gray!80blackDINOv2: vitb14 Pooling ERM (linear) 87.27 95.45 97.66 94.67 93.76
DI-ERM (linear) 87.35 96.53 98.05 94.50 94.11
gray!80blackDINOv2: vitl14 Pooling ERM (linear) 92.29 96.41 98.14 97.48 96.08
DI-ERM (linear) 92.42 97.37 98.10 97.48 96.34
DINOv3: vits16 Pooling ERM (linear) 85.14 94.25 94.34 92.92 91.67
DI-ERM (linear) 86.51 97.49 96.04 93.34 93.3
gray!80blackDINOv3: vitb16 Pooling ERM (linear) 93.61 79.58 94.58 97.53 91.33
DI-ERM (linear) 93.36 90.78 94.82 97.14 94.03
gray!80blackDINOv3: vitl16 Pooling ERM (linear) 96.18 92.28 94.04 99.1 95.40
DI-ERM (linear) 95.80 95.75 94.14 99.1 96.20